# Rule-based Data Selection for Large Language Models

### v2

@@ A.5.6 Prompts and Generated Rules
Comparison prompt: Below is the template used to compare two data samples according to a specific rule. For the rule-free version, simply omit the sentence involving the rule. Replace DATASET_NAME with “IMDB reviews” or “Common Crawl data” to correspond to the two data sources discussed in Section 4.
Compare two data examples from <DATASET_NAME> and choose the example which has better quality according to the following rule: <RULE>
The texts might have similar quality, but you should still make a relative judgement and choose the label of the preferred text. Example A: <DATA_SAMPLE A>
Example B: <DATA_SAMPLE B>
Now you have to choose between either A or B. You must respond only with a single letter ‘A’ or ‘B’.
@@
Index Rule and Description
@@
Table 5: Generated 50 Rules for rating IMDB examples.
@@
Index Rule Description
@@
Table 6: Generated 50 Rules for rating CommonCrawl examples.
A.6 Appendix for Evaluation B
A.6.1 Model training
@@
LoRA No Yes (rank=64)
A.6.2 More Baseline Methods
Here we add two more baseline methods: LESS (Xia et al., 2024): selecting data based on the estimated data influences, and DiverseEvol (Wu et al., 2023): an iterative sampling algorithm to ensure data diversity. It is important to note that DiverseEvol focuses solely on a single quality aspect: the diversity of data, while our method ensures diversity across multiple rating aspects. Another remark is that in the original papers, these methods were specifically used for instruction tuning data. We copy the three rule-based methods in our framework from 1 for comparison purposes. From the results below, we see these two methods, while being computationally expensive, are not showing good performance under our experiment settings.
Method ARC-Easy ARC-Challenge Winogrande MMLU SST-2 Average
LESS 59.8 25.2 52.9 25.6 49.3 42.5
DiverseEvol 59.7 24.7 53.3 25.6 49.2 42.5
All 50 Rules 60 25.3 53 26.1 49.3 42.7
Random 10 Rules 60 25.1 53.7 25.8 49.5 42.8
DPP 10 Rules 60 25.7 54 26.2 50.1 43.2
Table 8: General continued pre-training of Pythia-1B using 20K selected data samples from SlimPajama.
Method IMDB Medical
LESS 46.6 23.6 40.4 24 29.3
DiverseEvol 51.1 23.6 42.5 23 29.7
All 50 Rules 51 23.1 42.6 23 29.6
Random 10 Rules 51.7 24 41.2 23.5 29.6
DPP 10 Rules 53.5 24.6 43.3 26.8 31.6
Table 9: IMDB & Medical fine-tuning on Pythia-1B, each using 20K selected data samples from SlimPajama.
Method Math Code
LESS 41.5 40.4 33 38.3 41.4 43.5 43.9 45.3 43.5
DiverseEvol 41.2 38.5 35 38.2 38.4 43.6 42.8 47.8 43.1
All 50 Rules 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.8 35.2 39.3 48.5 41 46.6 48.1 46
DPP 10 Rules 43.7 40.6 38 40.8 50.5 44.2 46.9 52.7 48.6
Table 10: Math & Code fine-tuning on Llama3-8B, each using 20K selected data samples from SlimPajama.
A.6.3 Variance of Trials
Due to computational resource constraints, we were unable to perform multiple repetitions of all experiments. However, as mentioned in Section 5, we conducted 3 independent trials in four domains for Uniform Sampling and methods involving randomness in rule selections, including GPT-Uncorrelated, Random 10 Rules, and DPP 10 Rules (note that DPP sampling is also non-deterministic) to mitigate the effects of randomness, and we report their standard deviations here.
@@ A.6.3 Variance of Trials
Here we perform the $t$-test to demonstrate that the advantage of DPP 10 Rules is significant compared to other methods. We include the $t$-statistics and $p$-values in the table. If we choose the significance threshold $p=0.05$, then we see that all the comparisons are significant.
@@ A.6.4 Evaluation Benchmarks
In this section, we provide detailed descriptions of the benchmarks utilized for our evaluation. We considered the following benchmarks for general continued pre-training: ARC-Challenge (15), Winogrande (15), MMLU (5), SST-2 (0), where the numbers in parenthesis indicate the number of shots we use in few-shot benchmark setting. For domain fine-tuning, we use zero-shot in IMDB, and 5-shot for Medical and Math (which uses subsets of MMLU). Moreover, Math and Medical domains, we use the subject-related subsets from MMLU, specifically ElementaryMathematics, HighSchoolMathematics, and CollegeMathematics for Math, and CollegeMedicine, ProfessionalMedicine, and MedicalGenetics for Medical. For Code, we tested code generation and for each code benchmark, we use the pass@k setting and specify the number of code generation samples. See detailed explanations below.
  • MMLU (Maas et al., 2011): MMLU is a comprehensive multitask test comprises multiple-choice questions from a wide range of knowledge domains. It spans subjects across the humanities, social sciences, hard sciences, and other critical learning areas, encompassing 57 tasks such as elementary mathematics, US history, computer science, law, and more. To achieve high accuracy on this test, models need to demonstrate extensive world knowledge and robust problem-solving capabilities.
@@
  • IMDB (Maas et al., 2011): The IMDB dataset comprises 50,000 movie reviews and is designed for binary sentiment classification. For our evaluation, we select 25,000 test samples.
  • Winogrande (Sakaguchi et al., 2021): WinoGrande is a collection of 44,000 problems inspired by the Winograd Schema Challenge. It has been adjusted to enhance scale and robustness against dataset-specific bias. Designed as a fill-in-the-blank task with binary options, WinoGrande requires users to select the correct option for a given sentence based on commonsense reasoning.
  • SST-2 (Socher et al., 2013): SST-2, or the Stanford Sentiment Treebank binary classification dataset, is a widely used resource for sentiment analysis tasks. Derived from movie reviews, it consists of 11,855 single sentences, each annotated for sentiment polarity.
  • ARC-Easy and ARC-Challenge (Yadav et al., 2019): The AI2’s Reasoning Challenge (ARC) dataset is designed for evaluating multiple-choice question-answering systems. It consists of science exam questions for grades 3 to 9 and is divided into two subsets: Easy and Challenge. The Challenge subset comprises more complex questions that necessitate advanced reasoning skills. Typically, questions offer four answer choices, although a small fraction (less than 1%) may present three or five options. The dataset also features a Knowledge Base (KB) containing 14.3 million unstructured text passages to support reasoning and answer generation.
@@
  • HumanEval (Chen et al., 2021): The HumanEval benchmark evaluates Python programming skills with 164 problems, each comprising a function signature, docstring, function body, and unit tests. In a zero-shot setting, models generate code using top-p sampling (p=0.95) until stop words are reached. Pass@k metrics (k=1, 10, 100) are calculated with n=200 samples per problem, estimating the success rate following Chen et al.’s approach. Success is determined by whether at least one solution is correct within k attempts, with temperature controlling randomness in generation. This benchmark measures model performance in solving programming tasks with increasing attempts.
@@
  • MBPP (Austin et al., 2021): The MBPP benchmark contains around 1,000 crowd-sourced Python programming problems, designed for entry-level programmers. Each problem includes a task description, a code solution, and 3 test cases. The evaluation is performed on the test set from index 11 to 511. In a few-shot setting, the InCoder-style prompt is used, where the task description and one solution are provided to guide the model. The prompt format is f’"""{description}{test_example}"""’. By default, prompt_type_mbpp is set to incoder, and optionally, the solution can be included using include_solution_mbpp=True. We use single generation per problem (pass@1), and for pass@k estimation, we generate n=15 samples per problem, similar to the HumanEval approach. The evaluation focuses on pass@1 success rates.
@@
  • Multiple-py and Multiple-cpp (Cassano et al., 2022): MultiPL-E: is a benchmark for evaluating large language models for code generation that supports 18 programming languages. It takes the OpenAI “HumanEval” Python benchmark and uses little compilers to translate them to other languages. We use similar implementation as the original repository and evaluation parameters are similar to HumanEval.
A.6.5 Number of selected rules
@@
We modified the number of rules, $r$, from 10 to 20 and repeated the experiments for the Code domain. Compared to the 10-rule results presented in Table 3, we observed some discrepancies. For instance, the performance score on HumanEval is less than the 10-rule results, whereas the results for Multiple-cpp improved. The number of rules indeed alters the criteria used for data selection, thereby influencing the distribution of the selected data. Determining the optimal $r$ represents a valuable direction for future exploration.
@@
A.6.6 Size of Sampled Data
We investigated the impact of varying training data sizes on performance, specifically within the context of the DPP 10 rules and the Medical domain. Our observations reveal that increasing the amount of training data does not always enhance performance; in fact, performance may decline beyond a certain data threshold. This phenomenon is consistent with findings from the LIMA paper (Zhou et al., 2024), which suggests that data quality is often more important than quantity for LLMs. Balancing data quality with quantity is another challenging but valuable topic.
@@ A.6.6 Size of Sampled Data
A.6.7 Distribution of Selected Data
Evaluating and contrasting the quality of data subsets selected by different methods is challenging and often necessitates extensive human intervention. To address this, the authors examined the initial 100 examples selected by each method. This examination revealed notable distinctions in the relevance and domain specificity of the data selected. Specifically, our DPP rule-based approach demonstrated a marked ability to identify and select examples that were highly pertinent to specific domains. For instance, in experiments focused on the Code domain, this method favored the inclusion of data containing code. In contrast, other less targeted methods, such as QuRating and Uniform Sampling, often yield selections that lack domain-specific relevance. This insight underscores the efficacy of using tailored, rule-based methods over generic ones for tasks where domain alignment is critical.
@@ A.6.7 Distribution of Selected Data
Moreover for IMDB domain, in Figure 11 we investigated the text length distribution. We see that the QuRating is very close to the original SlimiPajama distribution, where we conjecture that in this case the data distribution is very close to uniformly sampled data. The methods within our framework have a tendency toward longer texts. Additionally, in Figure 12 we use bigram entropy (the Shannon entropy of the distribution over the unique bigrams) as an indicator of the text diversity. We again see that the entropy distribution of QuRating is very close to the original SlimPajama, where our methods generally select data with higher entropy/diversity and the entropy distributions are more concentrated.
@@
A.6.8 Rule Correlation of selected rules
Here we provide in Table 17 the rule indices (in the range $\{0,1,\dots,49\}$) for the rules selected by DPP and random selection (in one trial). The fullest of all generated $50$ rules for each domain is provided in A.6.11. For each set of selected rules, we also calculate their rule correlation value $\rho$ (defined in 1). We confirm that indeed DPP selects rules with lower rule correlation than random selected rules.
Table 17: Rule correlation and indices of the selected rules by DPP and random selection.
Domain Method Rule Indices
IMDB DPP 10 rules ($\rho=0.42$) [2, 3, 13, 21, 28, 36, 37, 45, 46, 49]
Random 10 rules A ($\rho=0.53$) [2, 6, 10, 11, 15, 21, 28, 42, 43, 48]
Random 10 rules B ($\rho=0.51$) [1, 9, 12, 14, 25, 26, 27, 37, 38, 40]
Medical DPP 10 rules ($\rho=0.55$) [1, 9, 10, 25, 29, 30, 32, 38, 42, 47]
Random 10 rules A ($\rho=0.69$) [11, 13, 14, 16, 25, 33, 34, 43, 45, 49]
Random 10 rules B ($\rho=0.66$) [6, 17, 20, 28, 29, 37, 40, 41, 47, 48]
Math DPP 10 rules ($\rho=0.40$) [0, 4, 13, 26, 27, 31, 33, 38, 44, 45]
Random 10 rules A ($\rho=0.65$) [0, 2, 11, 16, 17, 18, 27, 28, 34, 39]
Random 10 rules B ($\rho=0.61$) [3, 4, 25, 20, 23, 13, 15, 24, 35, 39]
Code DPP 10 rules ($\rho=0.54$) [2, 3, 13, 21, 28, 36, 37, 45, 46, 49]
Random 10 rules A ($\rho=0.59$) [5, 7, 10, 13, 17, 19, 21, 26, 30, 34]
Random 10 rules B ($\rho=0.58$) [2, 4, 8, 14, 16, 20, 23, 33, 37, 44]
A.6.9 GPT 10 uncorrelated rules
Another straightforward rule generation method is to directly prompt GPT-4 to generate 10 uncorrelated rules and rely on its understanding of the correlation between the rules. We have explored this by using a similar rule generation prompt as in A.6.11, where we provide the task description and data description, but this time we request for 10 rules and added one sentence “make sure the rules are uncorrelated” to further require the independence of the rules. The 10 “uncorrelated” GPT rules are provided in Table LABEL:tab:Appendix-EvalB-GPT10UncorrelatedRules-Code and LABEL:tab:Appendix-EvalB-GPT10UncorrelatedRules-Math below. Following this, we rated the data according to the 10 GPT rules and calculated the rule correlation $\rho$ of the score vectors (in one GPT-Uncorrelated trial). We tested for Code and Math domain and got $\rho_{Code}=0.65$ and $\rho_{Math}=0.56$, both are significantly higher than DPP correlation values in Tabel 17. For Code, even random 10 rules selected from a pool of 50 rules provide lower correlation than the 10 rules directly generated by GPT that are claimed to be “uncorrelated”. This shows that our two-step approach—first generating enough rules to ensure diversity, followed by employing DPP on the rating vectors to select rules—is superior and also more task-specific.
Index Rule Description
Index Rule Description
A.6.10 Use GPT to select 10 uncorrelated rules
In this part, we discuss a very similar setting to the previous section. However, instead of directly prompting GPT to generate 10 rules, we let GPT to replace the role of DPP and select 10 “uncorrelated” rules out of the 50-rule pool. First, in Table 20 below, we calculate the rule correlation similarly as in Table 17. We see again although we prompt GPT-4 to select “uncorrelated” rules, the rule-correlation of the selected 10 rules are still higher than our DPP-selected rules in Table 17. Moreover, we fine-tuned with the selected data and benchmarked the LLM performance. From the results in Table 21 and Table 22, we again see that it underperforms compared to our method.
Domain Method Rule Indices
IMDB GPT selected 10 rules ($\rho=0.67$) [0, 1, 4, 10, 13, 17, 25, 31, 40, 49]
Medical GPT selected 10 rules ($\rho=0.40$) [0, 4, 7, 11, 15, 24, 29, 34, 42, 49]
Math GPT selected 10 rules ($\rho=0.56$) [0, 3, 7, 11, 17, 24, 28, 38, 44, 48]
Code GPT selected 10 rules ($\rho=0.65$) [0, 4, 9, 12, 16, 23, 29, 34, 43, 49]
Table 20: Rule correlation and indices of the selected rules by DPP and random selection.
Method IMDB Medical
GPT selected 10 rules 51.6 21.9 42.2 24 29.4
All 50 Rules 51 23.1 42.6 23 29.6
Random 10 Rules 51.7 24 41.2 23.5 29.6
DPP 10 Rules 53.3 24.3 43 26 31.1
Table 21: IMDB & Medical fine-tuning on Pythia-1B, each using 20K selected data samples from SlimPajama.
Method Math Code
GPT selected 10 rules 42.5 40.3 36 39.6 40.8 43.4 42.8 50.3 44.3
All 50 Rules 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.6 35 39.2 48.5 40.8 46.6 48.1 46
DPP 10 Rules 43.6 40.4 38 40.7 50.6 44.1 46.9 52.8 48.6
Table 22: Math & Code fine-tuning on Llama3-8B, each using 20K selected data samples from SlimPajama.
A.6.11 Prompts and Generated Rules
@@
Generated 50 rules for each of four domains: Note that the IMDB rules here are used to select data for LLM training, whereas the IMDB rules in A.5.6 are used for data comparison in order to eventually calculate quality scores for the $50$ IMDB reviews. Although similar, they are not the same set of rules.
@@
Table 24: Generated 50 rules for the IMDB domain.
Index Rule Description
@@
Table 25: Generated 50 rules for the Medical domain.
Index Rule Description
@@
Table 26: Generated 50 rules for the Math domain.
Index Rule Description
@@
Table 27: Generated 50 rules for the Code domain.

### v3

@@ H.5 Prompts and Generated Rules
Below are the templates used to rate a sample according to a specific rule. For the rule-free version, simply omit the sentence involving the rule. Replace DATASET_NAME_AND_DESCRIPTION with the dataset’s name and one sentence of description to correspond to the two data sources discussed in Section 4.
Evaluate the following example from <DATASET_NAME_AND_DESCRIPTION> and assign a quality score between 0 and 1 (0 indicates the worst quality, and 1 indicates perfect quality) according to the provided rule: <RULE>
Example:
<DATA_SAMPLE>
Respond only with a single float number.
@@
The rules are then manually checked by 5 human annotations (all five are PhD students in CS or Math) to make sure the meaning is clear and not too abstract, and to remove the rules that are completely unrelated to the task/domain.
@@
Table 6: Generated rules for the IMDB examples.
@@
Table 7: Generated 50 Rules for rating CommonCrawl examples.
@@
Appendix I Appendix for Main Experiments
I.1 Model training
@@
LoRA No Rank=64
I.2 GPT 10 uncorrelated rules
Another straightforward rule generation method is to directly prompt GPT-4 to generate 10 uncorrelated rules and rely on its understanding of the correlation between the rules. We have explored this by using a similar rule generation prompt as in I.8, where we provide the task description and data description, but this time we request for 10 rules and added one sentence “make sure the rules are uncorrelated” to further require the independence of the rules. The 10 “uncorrelated” GPT rules are provided in Table 9 and 10 below. Following this, we rated the data according to the 10 GPT rules and calculated the rule correlation $\rho$ of the score vectors (in one GPT-Uncorrelated trial). We tested for Code and Math domain and got $\rho_{Code}=0.65$ and $\rho_{Math}=0.56$, both are significantly higher than DPP correlation values in Tabel 3. For Code, even random 10 rules selected from a pool of 50 rules provide lower correlation than the 10 rules directly generated by GPT that are claimed to be “uncorrelated”. This shows that our two-step approach—first generating enough rules to ensure diversity, followed by employing DPP on the rating vectors to select rules—is superior and also more task-specific.
Index Rule Description
Table 10: Generated 10 “uncorrelated” rules by GPT-4 for the Math domain.
Index Rule Description
I.3 Variance of Trials
Due to computational resource constraints, we were unable to perform multiple repetitions of all experiments. However, as mentioned in Section 5, we conducted 3 independent trials in four domains for Uniform Sampling and methods involving randomness in rule selections, including GPT-Uncorrelated, Random 10 Rules, and DPP 10 Rules (note that DPP sampling is also non-deterministic) to mitigate the effects of randomness, and we report their standard deviations here in Tables 11 and 12
@@ I.3 Variance of Trials
Here in Tables 13 and 14, we perform the $t$-test to demonstrate that the advantage of DPP 10 Rules is significant compared to other methods. We include the $t$-statistics and $p$-values in the table. If we choose the significance threshold $p=0.05$, then we see that all the comparisons are significant.
@@ I.4 Evaluation Benchmarks
In this section, we provide detailed descriptions of the benchmarks utilized for our evaluation. We considered the following benchmarks for fine-tuning: we use zero-shot in IMDB, and 5-shot for Medical and Math (which uses subsets of MMLU). Moreover, Math and Medical domains, we use the subject-related subsets from MMLU, specifically ElementaryMathematics, HighSchoolMathematics, and CollegeMathematics for Math, and CollegeMedicine, ProfessionalMedicine, and MedicalGenetics for Medical. For Code, we tested code generation and for each code benchmark, we use the pass@k setting and specify the number of code generation samples. See detailed explanations below.
@@
  • IMDB (maas-EtAl:2011:ACL-HLT2011): The IMDB dataset comprises 50,000 movie reviews and is designed for binary sentiment classification. For our evaluation, we select 25,000 test samples.
  • MMLU (maas-EtAl:2011:ACL-HLT2011): MMLU is a comprehensive multitask test comprises multiple-choice questions from a wide range of knowledge domains. It spans subjects across the humanities, social sciences, hard sciences, and other critical learning areas, encompassing 57 tasks such as elementary mathematics, US history, computer science, law, and more. To achieve high accuracy on this test, models need to demonstrate extensive world knowledge and robust problem-solving capabilities.
@@
  • HumanEval (chen2021evaluating): The HumanEval benchmark evaluates Python programming skills with 164 problems, each comprising a function signature, docstring, function body, and unit tests. In a zero-shot setting, models generate code using top-p sampling (p=0.95) until stop words are reached. Pass@k metrics (k=1, 10, 100) are calculated with n=200 samples per problem, estimating the success rate following Chen et al.’s approach. Success is determined by whether at least one solution is correct within k attempts, with temperature controlling randomness in generation. This benchmark measures model performance in solving programming tasks with increasing attempts.
@@
  • MBPP (austin2021program): The MBPP benchmark contains around 1,000 crowd-sourced Python programming problems, designed for entry-level programmers. Each problem includes a task description, a code solution, and 3 test cases. The evaluation is performed on the test set from index 11 to 511. In a few-shot setting, the InCoder-style prompt is used, where the task description and one solution are provided to guide the model. The prompt format is f’"""{description}{test_example}"""’.. By default, prompt_type_mbpp is set to incoder, and optionally, the solution can be included using include_solution_mbpp=True. We use single generation per problem (pass@1), and for pass@k estimation, we generate n=15 samples per problem, similar to the HumanEval approach. The evaluation focuses on pass@1 success rates.
@@
  • Multiple-py and Multiple-cpp (cassano2022multipl): MultiPL-E: is a benchmark for evaluating large language models for code generation that supports 18 programming languages. It takes the OpenAI “HumanEval” Python benchmark and uses little compilers to translate them to other languages. We use similar implementation as the original repository and evaluation parameters are similar to HumanEval.
I.5 Number of selected rules
The exploration of the scaling of $r$ in Section 4 is feasible due to the low cost of these experiments. However, in Section 5, which involves high-cost LLM training, it is impractical to test multiple values of $r$ across all experiments.
@@ I.5 Number of selected rules
Here in Table 15, we modified the number of rules, $r$, from 10 to 20 and repeated the experiments for the Code domain. Compared to the 10-rule results presented in Table 1, we observed some discrepancies. For instance, the performance score on HumanEval is less than the 10-rule results, whereas the results for Multiple-cpp improved. The number of rules indeed alters the criteria used for data selection, thereby influencing the distribution of the selected data. Determining the optimal $r$ represents a valuable direction for our future exploration.
@@
I.6 Size of Sampled Data
We investigated the impact of varying training data sizes on performance in Table 16, specifically within the context of the DPP 10 Rules and the Medical domain, due to limited computing resources. Our observations reveal that increasing the amount of training data does not always enhance performance; in fact, performance may decline beyond a certain data threshold. This phenomenon is consistent with findings from the LIMA paper (zhou2024lima), which suggests that data quality is often more important than quantity for LLMs. Balancing data quality with quantity is another challenging but valuable topic.
@@ I.6 Size of Sampled Data
I.7 Distribution of Selected Data
Evaluating and contrasting the quality of data subsets selected by different methods is challenging and often necessitates extensive human intervention. To address this, we examined the initial 100 examples selected by each method. This examination revealed notable distinctions in the relevance and domain specificity of the data selected. Specifically, our DPP rule-based approach demonstrated a marked ability to identify and select examples that were highly pertinent to specific domains. For instance, in experiments focused on the Code domain, this method favored the inclusion of data containing code. In contrast, other less targeted methods, such as QuRating and Uniform Sampling, often yield selections that lack domain-specific relevance. This insight underscores the efficacy of using tailored, rule-based methods over generic ones for tasks where domain alignment is critical.
@@ I.7 Distribution of Selected Data
Moreover for IMDB domain, in Figure 10 we investigated the text length distribution. We see that the QuRating is very close to the original SlimPajama distribution, where we conjecture that in this case the data distribution is very close to uniformly sampled data. The methods within our framework have a tendency toward longer texts. Additionally, in Figure 11 we use bigram entropy (the Shannon entropy of the distribution over the unique bigrams) as an indicator of the text diversity. We again see that the entropy distribution of QuRating is very close to the original SlimPajama, where our methods generally select data with higher entropy/diversity and the entropy distributions are more concentrated.
@@
I.8 Prompts and Generated Rules
@@
Generated 50 rules for each of four domains: Note that the IMDB rules here are used to select data for LLM training, whereas the IMDB rules in H.5 are used for data comparison in order to eventually calculate quality scores for the $50$ IMDB reviews. Although similar, they are not the same set of rules.
Table 18: Generated rules for the IMDB domain.
@@
Table 19: Generated rules for the Medical domain.
@@
Table 20: Generated rules for the Math domain.
@@
Table 21: Generated rules for the Code domain.
@@
Appendix H — Extended Discussion
We include here additional commentary on practical considerations, modeling decisions, and possible extensions of our method, with the aim of providing readers with a broader perspective on its applicability and trade-offs.
On the minimal manual filtering step.
While our pipeline is largely automated, we retain a light manual filtering step during rule generation to remove duplicate or malformed rules. This intervention is minimal: typically under 15 minutes per domain, and serves only to ensure syntactic clarity and diversity of candidate rules. In future applications, simple heuristics could easily replace this step. Importantly, once rule scoring and orthogonality-aware sampling begin, the selection process is entirely model-driven.
On compute requirements.
The cost of rule generation and rating is modest compared to full-scale fine-tuning. Rule generation uses a single GPT-4 call per domain, and the pairwise rule correlation matrix is computed efficiently over a single batch of n=10,000n=10{,}000 examples. The DPP selection step is fast due to the small size of the kernel ($R\leq 50$), and the entire pipeline runs comfortably on a single GPU. Overall, our method introduces only a minor overhead relative to the total training cost, while enabling more targeted, higher-quality supervision.
On task coverage.
We chose four representative domains: IMDB, Medical, Math, and Code to demonstrate the method’s generality across classification, reasoning, and program synthesis. These already cover a wide range of input modalities and supervision structures. Extending to more open-ended tasks (e.g. dialog safety or creative writing) is an exciting direction but would require new evaluation methodologies, which we view as orthogonal to the core contributions of this paper.
On rating model overlap.
Although we use the same backbone family (e.g., Llama-3) for scoring and fine-tuning, we mitigate circularity in two ways: (i) using both 8B and 70B models for rule-based scoring to assess robustness, and (ii) validating that downstream improvements hold across architectures and model sizes. We also include experiments using Claude-Sonnet as a rule generator to verify that our pipeline is not overly sensitive to the initial LLM used. Evaluating with independent human or hybrid judges is left for future work.
On parameter choices.
We fix $r=10$ rules per task and use $n=10K$ examples for scoring based on empirical pilot studies. These values work reliably across all domains. Our ablation study shows the method is stable across a broad range of $r$, and performance does not degrade sharply unless $r$ is extremely low or high. While further tuning may yield marginal improvements, our results suggest that high-quality rule-based selection does not require careful hyperparameter search.